the speed of a boat in still water is 9 km/hr.if the boat goes 54 km downstream in 4hrs,find the speed of the stream?
Answer by oscargut(2103)

total sspeed = 54/4 = 13.5
speed of the current = 13.5 - 9 = 4.5
Answer: 4.5 km/hr
You can ask me more at: mthman@gmail.com
Thanks